From: Mon 1/6/09



  • ATLAS ADCoS (prodsys shifters) meeting (phone CERN, Xavier Espinal) - skipped, overlapping
  • ATLAS TAG developers meeting (phone CERN, David Malon)
  • OSG Software Tools - postponed
  • fixed version of JSON server (LFC)
  • OSG Production
    • report on survey: 6 answers so far
  • Campfire:


  • ATLAS DB meeting (phone CERN)
  • ATLAS facilities meeting (phone EsNet)
  • Panda troubleshooting:
    • Problems with TAG analysis job: emails with Philippe Mermod <>,
    • dq2-put and GUIDs explanation to Vikas Bansal <>
    • Email to BNL and ddm list about short URL in AOD.064191._00486.pool.root.1 D8E932A1-A539-DE11-9158-0030487CF3EA ad:e69fabe8 1390602978
  • Campfire


  • setup CVS for local projects (vdtweb, dssddm)
  • prepare "faculty" convocation for OSG admin meeting
  • fixing Wlcg and DDM in OSG pages



ITB activity (site coordination)

Bulleted list: activity done, problems encountered

OSG Software activity

  • Discussion, email with Arvind about RSV
  • RSV installation on uct2-dss and tests

-- MarcoMambelli - 01 Jun 2009
Topic revision: r2 - 09 Jun 2009, MarcoMambelli
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback